The temperature coefficient of resistivity of carbon is - 0.0005/C°.if the resistor is made of graphite has a resistance of 5 k"\Omega" at 20° find the resistance at 27°







1
Expert's answer
2022-03-03T14:59:38-0500
"R(t_2)=R(t_1)(1+\alpha(t_2-t_1))"

"R(27)=R(20)(1+\alpha(27-20))"

"R(27)=5000\:\Omega(1+(- 0.0005)*7)=4983\:\Omega"